Gilbert makes frequent calls from country A to country B. His calling plan costs​ $5.00 per month plus 7¢ per minute for calls made to a landline and 17¢ per minute for calls made to a wireless number. One month his bill was ​$55.50. If he talked for a total of 350 ​min, how many minutes were to a landline and how many minutes to a wireless​ number?
Answer by ikleyn(52786)

.
The given problem is equivalent to this system of 2 equations in 2 unknowns

 x +   y = 350,                 (1)     (counting minutes)
7x + 17y = 5550 - 500           (2)     (counting dollars and cents)


After substitution  y = 350 - x  from the first equation to the second equation, the system is reduced to a single equation for only ONE unknown


7x + 17*(350-x) = 5050,

7x + 17*350 - 17x = 5050,

-10x = 5050 - 17*350 = -900  ====>  x = 90.


Answer.  90 landline minutes and the rest 350-90 = 260 wireless minutes.

Solved.   //   On the way, you learned on how the Substitution method works.

--------------------
To see other similar solved problems,  look into the lessons
    - Oranges and grapefruits
    - Using systems of equations to solve problems on tickets
    - Three methods for solving standard (typical) problems on tickets
    - Using systems of equations to solve problems on shares
    - Using systems of equations to solve problems on investment
    - Two mechanics work on a car
    - The Robinson family and the Sanders family each used their sprinklers last summer
    - Roses and vilolets
    - Counting calories and grams of fat in combined food
    - HOW TO algebreze and solve this problem on 2 equations in 2 unknowns
    - One unusual problem to solve using system of two equations
    - Solving mentally word problems on two equations in two unknowns
    - Solving systems of non-linear equations by reducing to linear ones
    - System of equations helps to solve a problem for the Thanksgiving day
    - Using system of two equations to solve the problem for the day of April, 1

    - OVERVIEW of lessons on solving systems of two linear equations in two unknowns
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Systems of two linear equations in two unknowns".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.